A potter's wheel is rotating around a vertical axis through its center at a frequency of 2.0 rev/s . The wheel can be considered a uniform disk of mass 4.7 kg and diameter 0.32 m . The potter then throws a 2.9-kg chunk of clay, approximately shaped as a flat disk of radius 7.0 cm , onto the center of the rotating wheel. Part A What is the frequency of the wheel after the clay sticks to it? Ignore friction.
Answer #1

Moment of inertia of wheel, Iw = 0.5 MR^2 = 0.5 x 4.7 x 0.16^2 = 0.06 kg m^2

Moment of inertia of clay, Ic = 0.5 x 2.9 x 0.07^2 = 0.0071 kg m^2

Using law of conservation of angular momentum,

I1 ω1 = I2 ω2

0.06 x 2 = (0.06 + 0.0071) x w2

w2 = 1.79 rad/sec

Comment in case any doubt please rate my answer....